A slack rope walker balances on a rope tied between two trees. For a short time, he is not moving. His mass is 72 kg. The rope makes an angle θ=22 with the horizontal on either side.

What is the magnitude of the tension in the rope?
1
Expert's answer
2020-08-14T15:57:35-0400

"2T\cos(90°-22°)=mg\to T=\frac{mg}{2\cos68°}=\frac{72\cdot9.81}{2\cos68°}\approx 943(N)"